OpenStudy (anonymous):

What number should replace the question mark? 19648, 1728, 112, ?

myininaya (myininaya):

those numbers are huge

myininaya (myininaya):

lol

OpenStudy (anonymous):

i need answer

myininaya (myininaya):

yes we are thinking

OpenStudy (anonymous):

that's the spirit

OpenStudy (anonymous):

no answer

myininaya (myininaya):

you can answer yourself if you are gonna keep rushing me i said i was thinking

OpenStudy (anonymous):

i thought i was on to something - 1728 is a perfect cube but the others are not!!

OpenStudy (anonymous):

how it came

OpenStudy (anonymous):

this isnt an obvious one

hero (hero):

What have you been studying in class?

OpenStudy (zarkon):

2

myininaya (myininaya):

lol how did you get 2

OpenStudy (zarkon):

1*9*6*4*8=1728 1*7*2*8=112 1*1*2=2

myininaya (myininaya):

neat

myininaya (myininaya):

gj zarkon

hero (hero):

this guy is a beast

OpenStudy (anonymous):

brilliant

myininaya (myininaya):

hes like totally smart

myininaya (myininaya):

i wasn't looking for that type of pattern lol i was way off on the approach

hero (hero):

I don't think anybody was

OpenStudy (anonymous):

i tried adding the digits together but why didn't i think of multiplying!!